the sum of twice a first number and fives times a second is 97. if the second number is subtracted from fives times the first number the result is 40.find the numbers.
Answer by mananth(16949)

one number =x
other =y
2x+5y=97-----1
5x-y=40-----2
multiply (2) by 5
25x-5y=200
add this to (1)
27x=297
/27
x=11
plug x in (1)
2x+5y=97
22+5y=97
5y=75
/5
y=15
11,15 are the numbers